[Uses the Marginal Utility Approach] Anita consumes both pizza and Pepsi. The following tables show the amount of utility she obtains from different amounts of these two goods:


Pizza

Pepsi


Quantity

Utility

Quantity

Utility


4 slices

115

5 cans

63


5 slices

135

5 cans

75


6 slices

154

5 cans

86


7 slices

171

5 cans

96






73-80. Suppose Pepsi costs $0.50 per can, pizza costs $1 per slice, and Anita has $9 to spend on food and drink. What combination of pizza and Pepsi will maximize her utility?


1
Expert's answer
2021-03-02T07:55:23-0500

Utility is maximized where MU of pizza/Price of Pizza = MU of Pepsi/Price of Pepsi 

"=(135-115)\/1=20"

"=(96-86)\/0.5=20"

5 canes of pizza and 5 canes of pepsi